What are the two main points of Law of demand? Illustrate law of demand with schedule and graph.

(04) 

(b) Define Elasticity of demand. Calculate price elasticity of demand for the given data:

(04)

 

Case A

Case B

Price

80

100

Quantity

230

150

 


1
Expert's answer
2021-08-29T16:51:19-0400

Solution:

The law of demand states that other things remaining constant, the amount demanded increases with a fall in price and diminishes with a rise in price. The law of demand explains the inverse relationship between the quantity demanded of a product and its price.

The two main points of the law of demand are:

1.     Ability by consumers to pay.

2.     Willingness by consumers to pay.

The price and demand of a product are inversely related to each other. When the price of a product increases, its demand falls. On the other hand, when the price of a product decreases, its demand rise.

The law of demand and schedule is illustrated below:

 


 



As per the above apples demand schedule and curve, we can observe that quantity demanded for apples is increasing with a fall in the price of apples. Hence, it is evident that the quantity demanded of a product is inversely related to its price.



b.). Elasticity of demand refers to an economic term that measures the change in the quantity demanded for a good or service relative to price changes for that particular good or service.

 

Price elasticity of demand (PEd) ="=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"


PEd = "\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"


Where: Q1 = 230, Q2 = 150, P1 = 80, P2 = 100


PEd = "\frac{150 -230}{(150+230)/2 }\div \frac{100 - 80}{(100+80)/2 }"


PEd = "\frac{-80}{190} \div\frac{20}{90}= \frac{-0.42}{0.22} = -1.91"


Price elasticity of demand (PEd) = 1.91